OpenStudy (anonymous):

What is the slope of the line that passes through the points (2,5) and (7,3)? Find the slope of the line passing through (11, -2) and (2, -2). Find the slope of the line passing through the points (-2, 9) and (5, 3). Show me how?

OpenStudy (anonymous):

formula of making linear equation is (y-y1)/(y2-y1)=(x-x1)/(x2-x1) at last the ratio of x is the slope

OpenStudy (anonymous):

1. -2/5 2. 0 3. -6/7 EXP: 1. (3-5)/(7-2)=-2/5 2. (-2--2)/(2-11)=0 3. (3-9)/(5--2) The formula is: (y2-y1)/(x2-x1)

OpenStudy (anonymous):

The rule is : y2-y1/x2-x1 just substitute the numbers in the question in the rule and you'll have the correct answer

OpenStudy (anonymous):

for example the first one will be : 3-5/7-2 = -2/5

OpenStudy (anonymous):

so i tried what you said on this Find the slope of line passing through the points (4,8) and (6, -2). and got -6/2 which is -3 but neither option is there.

OpenStudy (anonymous):

-2-8/6-4=-10/2=-5

OpenStudy (anonymous):

what did I do wrong?

OpenStudy (anonymous):

show me what u did ...maybe u mixed up the signs

OpenStudy (anonymous):

(4,8) and (6, -2) (2-8)/(6-4) -6/2

OpenStudy (anonymous):

u forgot to put the - sign infront of the 2

OpenStudy (anonymous):

ohhh Haha I get it now thanks!!!

OpenStudy (anonymous):

Happy to help :)

OpenStudy (anonymous):

okay so I feel really stupid for asking again... But I tried it on this one What is the slope of the line that passes through the points (–6,1) and (4,–4)? I didnt forget any - signs this time and

OpenStudy (anonymous):

i got -3/-2 and thats not an option

OpenStudy (anonymous):

-4-4/4-(-6)= -8/10=-4/5

OpenStudy (anonymous):

the two -- are turned into a +

OpenStudy (anonymous):

as u r writing it this way : -4-4/4+6=-8/10=-4/5

OpenStudy (anonymous):

as if *

OpenStudy (anonymous):

A.-2 B.2 C.-½ D.½ are my choices

OpenStudy (anonymous):

sorry ...my mistake i wrote -4 instead of -1

OpenStudy (anonymous):

here's the solution

OpenStudy (anonymous):

-4-1/4+6= -5/10=-1/2

OpenStudy (anonymous):

it's c .-1/2

OpenStudy (anonymous):

thanks I was sitting here going uhhh

OpenStudy (anonymous):

urw....again srry for the mixup :)